# GazeNoter: Co-Piloted AR Note-Taking via Gaze Selection of LLM Suggestions to Match Users’ Intentions

### v1

@@ Abstract.
datasets, neural networks, gaze detection, text tagging
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Embedded systemsccs: Computer systems organization Redundancyccs: Computer systems organization Roboticsccs: Networks Network reliability
@@ 1. Introduction
Note-taking is crucial during speech-based activities, such as in speeches and discussions. Besides summarizing content or organizing thoughts, it can act as real-time reminders for questions or opinions, particularly during question-and-answer (Q&A) sessions or when contributing timely to conversations. Common methods for note-taking include longhand notes and smartphone text input. However, manual note-taking can divert attention from the primary activity, distracting users and increasing their cognitive load (Piolat et al., 2005; Mosleh et al., 2016). This issue becomes more pronounced in mobile scenarios, such as taking notes during walking meetings. To address these challenges, prior research has utilized natural language processing (NLP) and large language models (LLMs) to automatically generate summaries (Xu et al., 2023; Shi et al., 2018), and highlights (Roy et al., 2021; Laban et al., 2023a) from a transcript context for note-taking. However, without user input, the auto-generated notes may not always align with the users’ intentions. To tackle this issue, recent studies (Kalnikaitundefined et al., 2012; Haliburton et al., 2023) have sought to involve user input in the automatic note-taking process by enabling users to highlight crucial statements during meetings. Nevertheless, these methods still fall short as they generate notes only within the context of the transcript but cannot produce notes or inferences that go beyond the context, which typically requires insights or inputs specific to the users. The importance of this type of derivative notes, which we refer to as beyond-context notes, has been emphasized by previous work in note-taking research (Fang et al., 2022).
@@
A nuanced system that incorporates user input into automatic note-taking, ensuring the alignment of generated content with users’ intentions, and providing support for both within-context and beyond-context notes is then required. Artificial intelligence (AI) and LLMs excel in diverse suggestion generation, and user input achieves more precise users’ intentions. By integrating user input into AI, forming a user-in-the-loop interaction paradigm, users can copilot with AI to adjust and generate more precise and desirable outcomes, which is leveraged in some works (Wang et al., 2022) but not for note-taking. Implementing this user-in-the-loop AI system for real-time note-taking requires a medium that facilitates swift, subtle, and low-distraction input and output manners for displaying and adjusting LLM-generated content to reduce users’ cognitive load and match users’ intentions. AR headsets with see-through displays, offering gaze selection, are ideal for this purpose, reducing the need to look down at smartphones and minimizing distraction (Cai et al., 2023; Janaka et al., 2022). Furthermore, AR interactions, enhanced by AI, can display content around real objects or humans, ensuring convenience and low distraction (Liu et al., 2023). Hence, an AR headset is proposed as the medium for this AI system, considering the expected rise in AR device usage in the future.
@@
We present an AI-copiloted augmented reality (AR) system, GazeNoter, enabling users to swiftly select LLM-generated suggestions via gaze on an AR headset for real-time note-taking. GazeNoter extracts keywords from the context of real-time audio transcripts via the LLM. Users can select these context keywords, or prompt the LLM to derive more keywords from the selected context keywords for further selection, aligning closely with users’ intentions. Using the selected keywords, the LLM then organizes candidate sentences that might encapsulate the desired notes for users to record the notes. By leveraging the three capabilities of the LLM, including extraction, derivation and organization, and efficient gaze selection for keyword and sentence selection, this AI-copilot AR system achieves real-time note-taking with low distraction and cognitive load, matching users’ intentions for both within-context and beyond-context notes. We conducted two user studies to compare the performance of GazeNoter with other methods, verifying its usability in attending speeches in a static sitting condition and walking meetings and discussions in a mobile walking condition, respectively.
@@
    The results from two 12-participant user studies, demonstrated GazeNoter’s effectiveness in static and dynamic scenarios over baseline comparisons, respectively. This effectiveness included reduced distraction and cognitive load, enhanced subtlety and usability, and improved resulting notes that better match users’ intentions and more effectively remind users. In the dynamic scenario, GazeNoter further improved frustration management, physical effort, and social acceptance.
@@ 2.2. Note-Taking Approaches
A variety of research has focused on facilitating the process of note-taking in lectures  (Kam et al., 2005; Silvestre et al., 2014; Plaue et al., 2012; Kaminski et al., 2016), meetings  (Damen et al., 2020a; Banerjee and Rudnicky, 2006, 2007; Exposito et al., 2017), online learning videos (Nguyen and Liu, 2016; Liu et al., 2019; Cao et al., 2022; Samson and Bassam, 2018) and other textual contents (Laban et al., 2023b; Rajaram and Nebeling, 2022). Assistive features are provided to enable more sophisticated interactions in note-taking systems. NoteLink (Srinivasa et al., 2021) retrieves relevant lecture videos based on students’ handwritten notes. NoteCoStruct (Fang et al., 2021) fosters a sense of learning community through note-sharing from previous learners. CoNotate (Palani et al., 2021) provides query suggestions based on the notes to assist exploratory searches in unfamiliar domains. However, text entry or handwriting is required in these works, which could divert users’ attention from ongoing activities, increasing cognitive load (Piolat et al., 2005; Mosleh et al., 2016). Besides, although voice recording is also used to take audio notes (Khan, 2019; Khan et al., 2022), which is proven more effective than typing notes using keyboard (Khan et al., 2022), broadcasting to listen to the notes is inappropriate in speech-based activities, such as speeches and discussions.
@@
On the other hand, several works focus on automatically generating summaries (Xu et al., 2023; Yang et al., 2022; Li et al., 2023) and highlights (Laban et al., 2023a; Hautasaari and Yamashita, 2014) to provide meaningful contents, especially with the development of LLMs. Li et al.  (Li et al., 2021) propose a system to generate different levels of summary and allow users to browse and navigate the content more efficiently. Beyond Text Generation (Dang et al., 2022) generates summaries to provide writers with an overview of their writing from an external perspective during the writing process. Note the Highlight (Roy et al., 2021) facilitates learning-oriented search tasks through an automatically highlighting search interface. TalkTraces (Chandrasegaran et al., 2019) introduces a real-time visualized summarization tool designed to assist in recognizing topics during discussions and gaining an overview of the agenda items discussed. MeetingVis (Shi et al., 2018) aims to facilitate productive group discussions during real-world problem-solving by presenting meeting elements through a visual, narrative approach. The system enhances meeting recall and encourages reflection on participants’ contributions while suggesting potential future research directions. These automatic note-taking methods are effective in producing summarized notes directly from the context of the transcripts, reducing cognitive load, but without users’ participation in the note-taking process, the notes within the context might not always align with those the users intend to record. Furthermore, they cannot generate beyond-context notes derived by users from the context, a factor whose importance has been shown in (Fang et al., 2022). On the other hand, The Walking Talking Stick (Haliburton et al., 2023) and Markup as you talk (Kalnikaitundefined et al., 2012) enable users to press a physical button to highlight critical statements during a meeting to discern and document users’ intentions. However, they might capture only segments of the transcript for within-context notes but not for beyond-context notes.
@@ 2.3. User-in-the-Loop NLP Systems
Instead of only using training models from computers in AI, previous research integrates an interactive process into AI systems and allows users to adjust and fine-tune the result with immediate feedback (Jin et al., 2017; Zhou and Yatani, 2022; Wu et al., 2022; Lu et al., 2023; Brade et al., 2023), which is also called the user-in-the-loop interaction paradigm. Such an interactive design enables the systems to generate more tailored, user-centered results. Moreover, the integration of user-in-the-loop interaction paradigm and with NLP has been effectively utilized across a variety of applications. ROPE (Wang et al., 2022) proposes an automatic audio shortening system based on the semantics and duration of the audio clip. It could optimize the result after users specify sentences to include or exclude. Crosspower (Xia, 2020) explores the utility of transforming linguistic structures extracted from written content into graphic content and enables users to directly interact with language structures and their graphic correspondences to create desired graphic effects. ConceptEVA (Zhang et al., 2023) proposes a document summarization system aimed at long and multi-domain documents to achieve the customization of long document summaries through an interactive visual analysis and NLP techniques.
Recent research has further incorporated user-in-the-loop with LLMs to improve the comprehension and generation capabilities of natural language in AI systems. Several research (Mirowski et al., 2023; Chung et al., 2022) offer users the flexibility to interact with the script generation process by LLMs and enrich the co-creation experience with LLMs in writing. Graphologue (Jiang et al., 2023) and Sensecape (Suh et al., 2023) have further introduced systems enabling users to interact with LLMs through non-linear, node-link dialogues. These systems allow users to dynamically tailor the graphical representation of information, offering flexibility in exploring and understanding complex knowledge. Moreover, Stylette (Kim et al., 2022a) utilizes an LLM to enable users to express their website stylistic preferences via natural language and generate a diverse palette of cascading style sheets (CSS) properties, offering users the flexibility to experiment and iterate to their desired website aesthetic. Promptify (Brade et al., 2023) introduces an interactive tool that enhances the prompt creation process for text-to-image models by offering a suggestion engine and a flexible interface for easy exploration and refinement. By integrating user input and AI, user-in-the-loop AI systems are able to generate more precise outcomes matching users’ intentions. Although these works do not focus on note-taking, we aim to leverage the merits of user-in-the-loop AI systems to enable a novel real-time AR note-taking system with low cognitive load, distraction, and matching users’ intentions for both within-context and beyond-context notes.
@@ 2.4. Gaze Selection and Layout on Headsets
Gaze is a common input method on VR/AR headsets. Rapid eye movement enables swift gaze selection, but performing confirmation using gaze is challenging. Previous research proposes several gaze-only target selection approaches, including dwelling (Pi and Shi, 2017; Thaler et al., 2013), gaze gestures (Dybdal et al., 2012), and eye vergence (Kudo et al., 2013; Kirst and Bulling, 2016; Ahn et al., 2020). Outline pursuits (Sidenmark et al., 2020) utilizes smooth pursuits of gaze movements to confirm selection among occluded objects. On the other hand, other works combine gaze and auxiliary modalities for gaze selection, including head movement (Sidenmark and Gellersen, 2019; Sidenmark et al., 2021; Wei et al., 2023), and hand gesture (Pfeuffer et al., 2017) on headsets. DEEP (Yi et al., 2022) leverages the eyelid movement to adjust the visual depth to select occluded targets. Gaze&Tongue (Gemicioglu et al., 2023) exploits subtle tongue gestures in gaze selection. Furthermore, some research focuses on designing a gaze-based menu to achieve more stable selections and reduce eye fatigue. StickyPie  (Ahn et al., 2021) and LatticeMenu (Kim et al., 2022b) focus on the design and layout of the marking menu for expert usage. Kuiper Belt (Choi et al., 2022) proposes placing menu items in an unnatural gaze angle to reduce false triggering.
Besides gaze selection, interacting with virtual elements via gaze on AR headsets without interrupting in-person conversations is another critical issue. StARe (Rivu et al., 2020) further presents a gaze-assisted interface to progressively reveal information about individuals on demand during a conversation. ParaGlassMenu (Cai et al., 2023) and Paracentral and Near-Peripheral Visualizations (Janaka et al., 2022) concentrated on designing the layout on optical see-through head-mounted displays (OHMD). Their goal was to maintain undiverted attention while performing other tasks, such as controlling digital devices or displaying information, on the OHMD. Based on these prior works, GazeNoter leverages the swift gaze selection in AR note-taking with a proper layout for social acceptance and subtlety.
@@ 3.1. Design Considerations
  • DC1. Real-Time System. In speech-based activities, e.g., speeches or discussions, taking notes for critical information, the essence of discourses, questions or opinions should be in a short period for the upcoming interactive Q&A sessions or discussions. Therefore, a real-time note-taking system that enables users to quickly take notes and preserve their thoughts at the moment becomes crucial. This means that instead of typing or speaking complete sentences for taking notes, swift input and output manners, such as only a few selection steps, for the user-in-the-loop note-taking system are required.
  • DC2. Matching Users’ Intentions. Without handwriting or text entry input, it is challenging to obtain users’ intentions in note-taking. Although LLMs excel in summarization and highlighting for note-taking, the automatically generated content as within-context notes could be diverse and might not always match users’ intentions since user input is not in the process. This also means that these automatic note-taking methods cannot generate beyond-context notes derived by users from the context, which is important in note-taking (Fang et al., 2022). By integrating user input into AI and forming a user-in-the-loop AI system, users can adjust the diverse LLM-generated output to converge closely to their intentions, generating more precise notes matching users’ intentions. For this user-in-the-loop AI system, there is a trade-off between generating more precise notes matching users’ intentions and avoiding complicated user input for swift input and lower distraction and cognitive load.
  • DC3. Distraction and Cognitive Load. Since users have to stay engaged in the ongoing activity while taking notes, note-taking via handwriting or smartphone text entry input could distract them and increase their cognitive load. Furthermore, the increased cognitive load may result in individuals taking incomplete notes (Kiewra, 1989). Note-taking methods with input and output manners that are either eyes-free or allow users to keep their visual attention on or around the ongoing activity are required. Eyes-free output methods, such as audio output, might be slow for real-time systems and have privacy issues. On the other hand, an always-worn AR headset with a see-through display to rapidly switch between virtual and real worlds for output that does not require users to look down at smartphones, allowing users to keep their visual attention on or around the ongoing activity could be an adequate medium. However, user input steps and time through the medium, and displayed content and layout on the medium might also affect users, potentially leading to distraction and cognitive load. These factors should be considered carefully.
@@
  • DC5. Mobility. Note-taking is required not only in static scenarios, such as seating or even with a desk, but also in mobile scenarios, such as attending walking meetings (Haliburton et al., 2023; Damen et al., 2020a, b). Therefore a portable design, including light-weight devices and even hands-free interactions for mobility should be considered.

### v2

@@ Abstract.
note-taking, augmented reality, large language models, artificial intelligence, gaze input, wearable devices
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Natural language interfacesccs: Human-centered computing Text input
@@
Users swiftly select LLM-generated suggestions via gaze on an AR headset with a ring for real-time note-taking in GazeNoter. (a) An AR headset and a ring are worn for note-taking in a speech. (b) Extracted context keywords from the latest sentence of the speech. (c) Users could select context and customized keywords, and candidate sentences are then automatically generated. (d) Users could explore and select derivative keywords beyond the context of the speech and select a candidate sentence best matching the intentions to record as a note. (e) Users could review the recorded notes, which may be from three processes. (Blue) Normally a sentence is recorded as a note. (Green) If no candidate sentences match users’ intentions, users could record all selected keywords as a note. (Red) If users need to take a note hastily, users could select context (or also customized) keywords, skipping step (d), to record these as a quick note.
@@ 1. Introduction
Note-taking is crucial during speech-based activities, such as in speeches and discussions. Besides summarizing content or organizing thoughts, it can act as real-time reminders for questions or opinions, particularly during question-and-answer (Q&A) sessions or when contributing timely to conversations. Common methods for note-taking include longhand notes and smartphone text input. However, manual note-taking can divert attention from the primary activity, distracting users and increasing their cognitive load (Piolat et al., 2005; Mosleh et al., 2016). This issue becomes more pronounced in mobile scenarios, such as taking notes during walking meetings. To address these challenges, prior research has utilized natural language processing (NLP) and large language models (LLMs) to automatically generate summaries (Xu et al., 2023; Shi et al., 2018), and highlights (Roy et al., 2021; Laban et al., 2023a) from a transcript context for note-taking. However, without user input, the auto-generated notes may not always align with the users’ intentions. To tackle this issue, recent studies (Kalnikaitundefined et al., 2012; Haliburton et al., 2023) have sought to involve user input in the automatic note-taking process by enabling users to highlight crucial statements during meetings. Nevertheless, these methods still fall short as they generate notes only within the context of the transcript but cannot produce notes or inferences that go beyond the context, which typically requires insights or inputs specific to the users, as illustrated in Figure 2. The importance of this type of derivative notes, which we refer to as beyond-context notes, has been emphasized by previous work in note-taking research (Fang et al., 2022).
@@
A nuanced system that incorporates user input into automatic note-taking, ensuring the alignment of generated content with users’ intentions, and providing support for both within-context and beyond-context notes is then required (the red block in Figure 2). Artificial intelligence (AI) and LLMs excel in diverse suggestion generation, and user input achieves more precise users’ intentions. By integrating user input into AI, forming a user-in-the-loop interaction paradigm, users can copilot with AI to adjust and generate more precise and desirable outcomes, which is leveraged in some works (Wang et al., 2022) but not for note-taking. Implementing this user-in-the-loop AI system for real-time note-taking requires a medium that facilitates swift, subtle, and low-distraction input and output manners for displaying and adjusting LLM-generated content to reduce users’ cognitive load and match users’ intentions. Augmented reality (AR) headsets with see-through displays, offering gaze selection, are ideal for this purpose, reducing the need to look down at smartphones and minimizing distraction (Cai et al., 2023; Janaka et al., 2022). Furthermore, AR interactions, enhanced by AI, can display content around real objects or humans, ensuring convenience and low distraction (Liu et al., 2023). Hence, an AR headset is proposed as the medium for this AI system, considering the expected rise in AR device usage in the future.
@@
We present an AI-copiloted AR system, GazeNoter, enabling users to swiftly select LLM-generated suggestions via gaze on an AR headset for real-time note-taking. GazeNoter extracts keywords from the context of real-time audio transcripts via the LLM. Users can select these context keywords, or prompt the LLM to derive more keywords from the selected context keywords for further selection, aligning closely with users’ intentions. Using the selected keywords, the LLM then organizes candidate sentences that might encapsulate the desired notes for users to record the notes. By leveraging the three capabilities of the LLM, including extraction, derivation and organization, and efficient gaze selection for keyword and sentence selection, this AI-copilot AR system achieves real-time note-taking with low distraction and cognitive load, matching users’ intentions for both within-context and beyond-context notes. While individual components, including LLMs for note-taking and AR gaze interactions, have been explored before, GazeNoter uniquely integrates these for mobile, real-time note-taking without voice or typing input. We conducted two user studies to compare the performance of GazeNoter with other note-taking methods, respectively verifying its usability in two scenarios: (1) attending speeches in a static sitting condition, similar to lecture scenarios widely studied in prior research (Mosleh et al., 2016; Piolat et al., 2005) but with more time constraints for note-taking; and (2) walking meetings and discussions in a mobile walking condition, an emerging use case (Damen et al., 2020a, b; Haliburton et al., 2023) with additional challenges due to mental pressure and the walking environment.
@@
    The results from two user studies, demonstrated GazeNoter’s effectiveness in static and mobile scenarios over baseline comparisons, respectively. This effectiveness included reduced distraction and cognitive load, enhanced subtlety and usability, and improved resulting notes that better match users’ intentions and more effectively remind users. In the mobile scenario, GazeNoter further improved frustration management, physical effort, and social acceptance.
Figure 2. The blue part represents the notes users want to directly record as hearing the context, which is defined as within-context notes. The green part represents notes that combine users’ insights, defined as beyond-context notes. Achieving both within-context and beyond-context notes ensures that the notes align with users’ intentions.
[…]
The blue part represents the notes users want to directly record as hearing the context, which is defined as within-context notes. The green part represents notes that combine users’ insights, defined as beyond-context notes. Achieving both within-context and beyond-context notes ensures that the notes align with users’ intentions.
@@ 2.2. Note-Taking Approaches
A variety of research has focused on facilitating the process of note-taking in lectures  (Kam et al., 2005; Silvestre et al., 2014; Plaue et al., 2012; Kaminski et al., 2016), meetings  (Damen et al., 2020a; Banerjee and Rudnicky, 2006, 2007; Exposito et al., 2017), online learning videos (Nguyen and Liu, 2016; Liu et al., 2019; Cao et al., 2022; Samson and Bassam, 2018) and other textual contents (Laban et al., 2023b; Rajaram and Nebeling, 2022). Assistive features are provided to enable more sophisticated interactions in note-taking systems. NoteLink (Srinivasa et al., 2021) retrieves relevant lecture videos based on students’ handwritten notes. NoteCoStruct (Fang et al., 2021) fosters a sense of learning community through note-sharing from previous learners. CoNotate (Palani et al., 2021) provides query suggestions based on the notes to assist exploratory searches in unfamiliar domains. However, text entry or handwriting is required in these works, which could divert users’ attention from ongoing activities, increasing cognitive load (Piolat et al., 2005; Mosleh et al., 2016). Besides, although voice recording is also used to take audio notes (Khan, 2019; Khan et al., 2022), which is proven more effective than typing notes using keyboard (Khan et al., 2022), broadcasting to listen to the notes is inappropriate in activities like speeches and discussions.
@@
On the other hand, several works focus on automatically generating summaries (Xu et al., 2023; Yang et al., 2022; Li et al., 2023) and highlights (Laban et al., 2023a; Hautasaari and Yamashita, 2014) to provide meaningful contents, especially with the development of LLMs. TalkTraces  (Chandrasegaran et al., 2019)and MeetingVis (Shi et al., 2018) focus on real-time visualized summarization tools designed to facilitate productive group discussions by recognizing topics and providing a visualized overview of the agenda in real-time. Li et al.  (Li et al., 2021) propose a system to generate different levels of summary and allow users to browse and navigate the content more efficiently. Beyond Text Generation (Dang et al., 2022) generates summaries to provide writers with an overview of their writing from an external perspective during the writing process. These automatic note-taking methods are effective in producing summarized notes directly from the context of the transcripts, reducing cognitive load, but without users’ participation in the note-taking process, the notes within the context might not always align with those the users intend to record. Furthermore, they cannot generate beyond-context notes derived by users from the context, a factor whose importance has been shown in (Fang et al., 2022). On the other hand, The Walking Talking Stick (Haliburton et al., 2023) and Markup as you talk (Kalnikaitundefined et al., 2012) enable users to press a physical button to highlight critical statements during a meeting to discern and document users’ intentions. However, they might capture only segments of the transcript for within-context notes but not for beyond-context notes.
@@ 2.3. User-in-the-Loop NLP Systems
Instead of only using training models, previous research integrates an interactive process into AI systems and allows users to adjust and fine-tune the result with immediate feedback (Jin et al., 2017; Zhou and Yatani, 2022; Wu et al., 2022; Lu et al., 2023; Brade et al., 2023), which is also called the user-in-the-loop interaction paradigm. Such an interactive design enables the systems to generate more tailored, user-centered results. The integration of user-in-the-loop interaction paradigm and with NLP has been effectively utilized across a variety of applications. ROPE (Wang et al., 2022) proposes an automatic audio shortening system based on the semantics and duration of the audio clip and can optimize results after users specify sentences to include or exclude. Crosspower (Xia, 2020) explores the utility of transforming linguistic structures extracted from written content into graphic content and enables users to interact with language structures and their graphic correspondences to create graphic effects. ConceptEVA (Zhang et al., 2023) proposes a document summarization system to achieve the customization of long document summaries through an interactive visual analysis and NLP techniques.
Recent research has further incorporated user-in-the-loop with LLMs to improve the comprehension and generation capabilities of natural language in AI systems. Several research (Mirowski et al., 2023; Chung et al., 2022) offer users the flexibility to interact with the script generation process by LLMs and enrich the co-creation experience with LLMs in writing. Graphologue (Jiang et al., 2023) and Sensecape (Suh et al., 2023) propose systems that enable users to interact with LLMs through non-linear, node-link dialogues and allow users to dynamically tailor the graphical representation of information, offering flexibility in exploring and understanding complex knowledge. Promptify (Brade et al., 2023) introduces an interactive tool that enhances the prompt creation process for text-to-image models by offering a suggestion engine and a flexible interface for easy exploration and refinement. By integrating user input and AI, user-in-the-loop AI systems are able to generate more precise outcomes matching users’ intentions. Although these works do not focus on note-taking, we aim to leverage the merits to enable a novel real-time AR note-taking system with low cognitive load, distraction, and matching users’ intentions for both within-context and beyond-context notes.
@@ 2.4. Gaze Selection and Layout on Headsets
Gaze is a common input method on VR/AR headsets. Rapid eye movement enables swift gaze selection, but performing confirmation using gaze is challenging. Previous research proposes several gaze-only target selection approaches, including dwelling (Pi and Shi, 2017; Thaler et al., 2013), gaze gestures (Dybdal et al., 2012; Sidenmark et al., 2020), and eye vergence (Kudo et al., 2013; Kirst and Bulling, 2016; Ahn et al., 2020). On the other hand, others combine gaze and auxiliary modalities for gaze selection, including head movement (Sidenmark and Gellersen, 2019; Sidenmark et al., 2021; Wei et al., 2023), eyelid movement (Yi et al., 2022), tongue gesture (Gemicioglu et al., 2023) and hand gesture (Pfeuffer et al., 2017) on headsets. Furthermore, some research focuses on designing a gaze-based menu to achieve more stable selections and reduce eye fatigue (Ahn et al., 2021; Kim et al., 2022; Choi et al., 2022). In gaze selection, interacting with virtual elements via gaze on AR headsets without interrupting conversations is another critical issue. StARe (Rivu et al., 2020) and Glanceable AR (Lu et al., 2020) focus on designing progressively revealing information on demand to minimize distraction using gaze on headsets. ParaGlassMenu (Cai et al., 2023) and Paracentral and Near-Peripheral Visualizations (Janaka et al., 2022) concentrated on designing the layout on OHMDs. Their goal was to maintain undiverted attention while performing other tasks, such as controlling digital devices or displaying information, on the OHMD. Based on these prior works, GazeNoter leverages the swift gaze selection in AR note-taking with a proper layout for social acceptance and subtlety.
@@ 3.1. Design Considerations
  • DC1. Real-Time System. In speech-based activities, e.g., speeches or discussions, taking notes for critical information, the essence of discourses, questions or opinions should be in a short period for the upcoming interactive Q&A sessions or discussions. Therefore, a real-time note-taking system that enables users to quickly take notes and preserve their thoughts at the moment becomes crucial. This means that instead of typing or speaking complete sentences for taking notes, swift input and output manners, such as only a few selection steps, for the note-taking system are required.
  • DC2. Matching Users’ Intentions. Without handwriting or text entry input, it is challenging to obtain users’ intentions in note-taking. Although LLMs excel in summarization and highlighting for note-taking, the automatically generated content as within-context notes could be diverse and might not always match users’ intentions since user input is not in the process. This also means that these automatic note-taking methods cannot generate beyond-context notes derived by users from the context. Moreover, previous studies have underscored the critical role of notes that incorporate derivative words and go beyond the given context to improve the quality of notes. Such notes are characterized as being constructive and interactive within the adaptive Interactive, Constructive, Active, and Passive (ICAP) framework (Fang et al., 2022), which categorizes note quality to enhance the learning experience. By integrating user input into AI and forming a user-in-the-loop AI system, users can adjust the diverse LLM-generated output to converge closely to their intentions, generating more precise notes matching users’ intentions. For such a system, there is a trade-off between generating more precise notes matching users’ intentions and avoiding complicated user input for swift input and lower distraction and cognitive load.
  • DC3. Distraction and Cognitive Load. Since users have to stay engaged in the ongoing activity while taking notes, note-taking via handwriting or smartphone text entry input could distract them and increase their cognitive load. The increased cognitive load may result in individuals taking incomplete notes (Kiewra, 1989). Thus, note-taking methods with input and output manners that are either eyes-free or allow users to keep their visual attention on or around the ongoing activity are required. Eyes-free output methods, such as audio output, might be slow for real-time systems and have privacy issues. On the other hand, an always-worn AR headset with a see-through display to rapidly switch between virtual and real worlds for output that does not require users to look down at smartphones, allowing users to keep their visual attention on or around the ongoing activity could be an adequate medium. However, user input steps and time through the medium, and displayed content and layout on the medium might also affect users, potentially leading to distraction and cognitive load, which should be considered carefully.
@@
  • DC5. Mobility. Note-taking is required in static scenarios, such as seating or even with a desk, and in mobile scenarios, such as attending walking meetings (Haliburton et al., 2023; Damen et al., 2020a, b). Thus, a portable design, including lightweight devices and even hands-free interactions for mobility should be considered.
@@
(Middle) The flowchart of GazeNoter. (Right) The AR layout, which is shown only when the ring is touched. (a) GazeNoter extracts context keywords from the latest sentence of the speech. (b) Once the user selects a context keyword, ”city”, the pre-defined customized keywords are shown, and 3 candidate sentences are generated based on the context keyword. (c) The user selects a customized keyword, ”what”, and the candidate sentences are updated accordingly. (d) If the user wants to take a beyond-context note and no desired keyword is among the context keywords, the user selects the most relevant context keyword, ”rallies”, to generate derivative keywords. (e) The user selects a derivative keyword, ”sign”, and the candidate sentences beyond the context are updated accordingly. (f) The user selects the candidate sentence best matching the intention, ”What signs were displayed in each city?”, to record as a note. (g) The recorded note is shown. (h) If no candidate sentences match the intention in step (f), the user could also record all selected keywords as a note (upper). If the user needs to take a note hastily, users could select only context (or also customized) keywords to record these as a quick note (lower), only from steps (a)(b) or (a)(b)(c).